# Functional Clustering for Longitudinal Associations between Social Determinants of Health and Stroke Mortality in the US

### v2

@@ 2.2 Finite Mixture of Functional Linear Concurrent models
Let $Y_{i}(t)\in\mathbb{R}$ and $\bm{X}_{i\cdot}(t):=(X_{i1}(t),\dots,X_{ip}(t))^{T}\in\mathbb{R}^{p}$ represent the stroke mortality and SDOH data in the $i$th county at time $t$, respectively, where $i=1,\ldots,n$ and $t\in\mathcal{T}$, with $n$ being the number of counties and $\mathcal{T}$ being the observed time period. Here, $Y_{i}(\cdot)$ and $\bm{X}_{i\cdot}(\cdot)$ are considered the functional response and covariate samples, respectively, with $n$ and $p$ being the sample size and dimension of the covariates. Without loss of generality, $\mathcal{T}$ is $[0,1]$ in this article.
@@
We focus on the association between $Y_{i}(t)$ and $\bm{X}_{i\cdot(t)}$ for $t\in[0,1]$, referred to as longitudinal associations in what follows. To cluster the longitudinal associations of different counties $i$, we assume the samples $\{Y_{i}(\cdot),{\bm{X}_{i\cdot}(\cdot)};i=1,\ldots,n\}$ can be divided into $K$ clusters and follow a finite mixture of functional linear concurrent models
$Y_{i}(t)=\sum_{j=1}^{p}X_{ij}(t)\beta_{jk}(t)+\varepsilon_{ik}(t)\ \text{if}\ Z_{i}=k,$ (1)
where $Z_{i}$ is the cluster membership for the $i$th subject, $\beta_{jk}(\cdot)$ is the functional coefficient for the $k$th group to capture the longitudinal associations between the response and the $j$th covariate, and $\varepsilon_{ik}(t)$ is a Gaussian white noise with variance $\sigma_{k}^{2}$ for each $k$. We assume that the white noise processes $\{\varepsilon_{ik}(\cdot);i=1,\ldots,n,k=1,\ldots,K\}$ are independent across different $i$s and $k$s, and $Z_{1},\ldots,Z_{n}$ are i.i.d. samples from a multinomial distribution on $\{1,\ldots,K\}$ with $\mathbb{P}(Z_{i}=k)=\pi_{k}$.
@@
Let $\bm{\beta}(t)=\left\{\beta_{jk}(t)\right\}_{j=1,\ldots,p;k=1,\ldots,K}$, $\bm{\sigma}^{2}:=(\sigma_{1}^{2},\ldots,\sigma_{K}^{2})^{T}$ and $\bm{\pi}:=(\pi_{1},\ldots,\pi_{K})^{T}$. We use $\bm{\beta}_{\cdot k}(t)$ and $\bm{\beta}_{j\cdot}(t)$ to denote the $k$th column and the $j$th row of $\bm{\beta}(t)$, respectively. For ease of notation, we might use $\bm{\beta}$, $\bm{\beta}_{\cdot k}$, $\bm{\beta}_{j\cdot}$, and $\beta_{jk}$ to represent $\bm{\beta}(\cdot)$, $\bm{\beta}_{\cdot k}(\cdot)$, $\bm{\beta}_{j\cdot}(\cdot)$ and $\beta_{jk}(\cdot)$ in what follows.
Since our dataset are only observed on a finite time grid, we assume that $\big{\{}\big{(}Y_{i}(t_{is}),\bm{X}_{i}(t_{is})\big{)};$ $i=1,\ldots,n,s=1,\ldots,S_{i}\big{\}}$ is the observed data for FMFLCM, where $\{t_{is};i=1,\ldots,n,s=1,\ldots,S_{i}\}\subset\mathcal{T}$. The log-likelihood of the parameters ${\bm{\Phi}}:=(\bm{\beta},\bm{\sigma}^{2},\bm{\pi})$ is then given as
$l({\bm{\Phi}})$ $=\sum_{i=1}^{n}\sum_{s=1}^{S_{i}}\text{log}\left\{\sum_{k=1}^{K}{\pi_{k}}f\left(Y_{i}(t_{is});\bm{X}_{i}(t_{is})^{T}\bm{\beta}_{\cdot k}(t_{is}),\sigma_{k}^{2}\right)\right\},$ (2)
where $f(\cdot;\mu,\sigma^{2})$ is a Gaussian density function with mean $\mu$ and variance $\sigma^{2}$. One may adopt the maximum likelihood estimator (MLE) of (2) to estimate $\bm{\Phi}$. However, given that the functional coefficient $\bm{\beta}$ is an infinite-dimensional parameter, the optimization of $\bm{\beta}$ in (2) is generally an ill-posed problem. Furthermore, the MLE of $\bm{\beta}$ may also suffer from the curse of dimensionality, i.e., $p$ is large. For this case, the MLE of (2) may be far from its true value, or even does not exist (Wang et al., 2008; Yi and Caramanis, 2015).
To solve the above issues, we propose a class of penalties in Section 2.3 to regularize the smoothness and the sparsity of $\bm{\beta}$. Based on the proposed penalties, we develop a regularized expectation maximization algorithm in Section 2.4 to implement the clustering for subjects, the estimation for $\bm{\beta}$, and the variable selections for the covariates simultaneously.
@@ 2.3 Sparsity and Smoothness Pursued Penalties
In this subsection, we propose a class of smoothly clipped absolute deviation (SCAD) type penalties (Fan and Li, 2001) to adjust the aforementioned ill-posed problems. The SCAD penalty takes the form
$\bm{P}_{\text{SCAD}}\left(u;\lambda\right)=\left\{\begin{array}[]{ll}\lambda u&\text{if}\ 0\leq u\leq\lambda,\\ -\frac{(u^{2}-2\gamma\lambda u+\lambda^{2})}{2(\gamma-1)}&\text{if}\ \lambda<u<\gamma\lambda,\\ \frac{(\gamma+1)\lambda^{2}}{2}&\text{if}\ u\geq\gamma\lambda,\end{array}\right.$
@@
where $u$ is a scalar parameter to be penalized, $\lambda$ is a tuning parameter, and $\gamma$ is a hyperparameter chosen to be 3.7 as suggested in Fan and Li (2001). We extend the SCAD penalty for the functional objects $\bm{\beta}$, which is the functional SCAD (FSCAD) penalty
@@
$\bm{P}_{\text{FSCAD}}(\bm{\beta};{\lambda},r):=\sum_{j=1}^{p}\sum_{k=1}^{K}\bm{P}_{\text{SCAD}}(\|\beta_{jk}\|_{r};{\lambda}),$ (3)
@@
where $\|\beta_{jk}\|_{r}=\sqrt{\|\beta_{jk}\|^{2}+r\|\beta_{jk}^{{}^{\prime\prime}}\|}$ with $\|f\|=\sqrt{\int f^{2}(t)\ \mathrm{d}t}$ being an $L_{2}$ norm for a function $f$. Here, $\|\beta_{jk}\|_{r}$ measures the magnitude and smoothness of $\beta_{jk}$ simultaneously (Meier et al., 2009), in which $r$ leverages the function’s magnitude and smoothness in the norm. By penalizing $\bm{\beta}$ using (3), we can penalize the smoothness of each $\beta_{jk}$ to deal with its functional nature, and shrink the $\beta_{jk}$ with small norm $\|\beta_{jk}\|_{r}$ to zero (Huang et al., 2009) for the purpose of variable selections.
@@
It’s worth noting that our SDOH functional covariates exhibit significant collinearity as illustrated in Figure 1, and the presence of the collinearity may enlarge variance and lead to misspecification in variable selections (Zou and Hastie, 2005). To adjust for this issue, we adopt a similar strategy as the elastic net (Zou and Hastie, 2005) to add an $L_{2}$ term to the FSCAD penalty (3). As such, we obtain the functional SCAD-Net (FSCAD-Net) penalty
$\bm{P}_{\text{FSCAD-Net}}(\bm{\beta};\lambda,\rho,r):=\sum_{j=1}^{p}\sum_{k=1}^{K}\bigg{\{}\rho\bm{P}_{\text{FSCAD}}\left(\beta_{jk};\lambda,r\right)+(1-\rho){\color[rgb]{.75,.5,.25}{\lambda_{k}}}\|\beta_{jk}\|_{r}^{2}\bigg{\}},$ (4)
@@
where $\rho\in[0,1]$ is a tuning parameter controlling the proportions between the FSCAD and $L_{2}$ terms. The above penalty maintains the properties of the FSCAD penalty. In addition, the FSCAD-Net penalty can reduce variance and address misspecification in variable selections by penalizing the magnitude of $\sum_{j=1}^{p}\sum_{k=1}^{K}\|\beta_{jk}\|_{r}^{2}$.
@@
It’s worth noting that the variable selections with FSCAD or FSCAD-Net do not guarantee that the selected covariates from different clusters are the same. This may not be reasonable for some applications, in which we need to identify the important variables for all clusters. In our case, the purpose of variable selections for SDOH is to facilitate the clustering on the selected SDOH covariates. To this end, we require the selected SDOH for all clusters to be the same within the estimation procedure. To facilitate such kind of investigation, we propose to shrink $\{\beta_{jk};k=1,\cdots,K\}$ to zero functions simultaneously within the variable selection, which is referred to as cluster-invariant sparsity. For this, we modify the FSCAD-Net penalty into the functional group SCAD-Net (FGSCAD-Net) penalty
@@
$\bm{P}_{\text{FGSCAD-Net}}(\bm{\beta};\lambda,\rho,r):=\sum_{j=1}^{p}\bigg{\{}\rho\bm{P}_{\text{SCAD}}(\|\bm{\beta}_{j\cdot}\|_{r};\lambda)+(1-\rho)\lambda\|\bm{\beta}_{j\cdot}\|_{r}^{2}\bigg{\}},$ (5)
@@
where $\|\bm{\beta}_{j\cdot}\|_{r}=\sqrt{\sum_{k=1}^{K}\|\beta_{jk}\|_{r}^{2}}$. By grouping $\left\{\beta_{kj};k=1,\ldots,K\right\}$ for each $j$ together in the norm, the FGSCAD-Net not only yields cluster-invariant sparsity but also improves statistical efficiency for variable selection by borrowing strength from all clusters for estimating $\bm{\beta}$.
@@ 2.4 Regularized EM Algorithm
In this subsection, we apply an EM algorithm to conduct the parameter estimation of FMFLCM. We first introduce a latent variable $z_{ik}:=I(Z_{i}=k)$ to represent the cluster membership of the subject $i$, where $I(\cdot)$ is an indicator function. Denote $\bm{Z}$ as $\{z_{ik};i=1,\ldots,n,k=1,\ldots,K\}$. The complete log-likelihood of FMFLCM containing $\bm{Z}$ is then given by
$l_{C}(\bm{\Phi},\bm{Z})$ $=\sum_{i=1}^{n}\sum_{k=1}^{K}z_{ik}\sum_{s=1}^{S_{i}}\bigg{[}\text{log}\pi_{k}+\text{log}\left\{f\left(Y_{i}(t_{is});\{\bm{X}_{i}(t_{is})\}^{T}\bm{\beta}_{\cdot k}(t_{is}),\sigma_{k}^{2}\right)\right\}\bigg{]}.$ (6)
In order to solve the ill-posed issue for estimating $\bm{\beta}$, we employ a regularized EM (REM) algorithm using the FGSCAD-Net penalty (5) as proposed in Section 2.3. In the following, we demonstrate the E-step and M-step of the REM algorithm.
2.4.1 General procedures of REM algorithm
Let $\bm{\Phi}^{(m-1)}:=(\bm{\beta}^{(m-1)},(\bm{\sigma}^{2})^{(m-1)},\bm{\pi}^{(m-1)})$ be the parameters updated at the $(m-1)$th iteration.
E-step: With the parameter $\bm{\Phi}^{(m-1)}$, we first compute the conditional expectation of $z_{ik}$ given the samples $\bm{Y}=\{Y_{i}(t_{is});i=1,\ldots,n,s=1,\ldots.S_{i}\}$ and $\bm{X}=\{\bm{X}_{i}(t_{is});i=1,\ldots,n,s=1,\ldots,S_{i}\}$
$\omega_{ik}^{(m)}:=\mathbb{E}_{\bm{\Phi}^{(m-1)}}(z_{ik}\mid\bm{Y},\bm{X})=\frac{{\pi_{k}^{(m-1)}}\sum_{s=1}^{S_{i}}f\left(Y_{i}(t_{is}),\bm{X}_{i}(t_{is})^{T}\bm{\beta}_{\cdot k}^{(m-1)}(t_{is}),(\sigma_{k}^{2})^{(m-1)}\right)}{\sum_{k=1}^{K}{\pi_{k}^{(m-1)}}\sum_{s=1}^{S_{i}}f\left(Y_{i}(t_{is}),\bm{X}_{i}(t_{is})^{T}\bm{\beta}_{\cdot k}^{(m-1)}(t_{is}),(\sigma_{k}^{2})^{(m-1)}\right)}.$ (7)
After that, we calculate the expectation of the complete log-likelihood in (6) conditioning on $\bm{Y}$ and $\bm{X}$, given the parameter $\bm{\Phi}^{(m-1)}$. This leads to the Q function
$Q(\bm{\Phi}\mid\bm{\Phi}^{(m-1)}):=\sum_{i=1}^{n}\sum_{k=1}^{K}\omega_{ik}^{(m)}\sum_{s=1}^{S_{i}}\bigg{[}\text{log}\pi_{k}+\text{log}\left\{f\left(Y_{i}(t_{is}),\{\bm{X}_{i}(t_{is})\}^{T}\bm{\beta}_{\cdot k}(t_{is}),\sigma_{k}^{2}\right)\right\}\bigg{]}.$ (8)
M-step: To facilitate the update of $\bm{\beta}$, we incorporate the FGSCAD-Net penalty (5) into the Q function (8).
$Q^{\text{pen}}\left(\bm{\Phi}\mid\bm{\Phi}^{(m-1)};\lambda,\rho,r\right):=Q\left(\bm{\Phi}\mid\bm{\Phi}^{(m-1)}\right)-\bm{P}_{\text{FGSCAD-Net}}\left(\bm{\beta};\lambda,\rho,r\right).$ (9)

### v6

@@ 2.2 Finite Mixture of Functional Linear Concurrent Models
For $i=1,\ldots,n$ and $t\in\mathcal{T}$, let $Y_{i}(t)\in\mathbb{R}$ and $\bm{X}_{i}(t):=(X_{i1}(t),\dots,X_{ip}(t))^{\top}\in\mathbb{R}^{p}$ represent the stroke mortality and SDOH data in the $i$th county at time $t$, respectively, where $n$ is the number of counties and $\mathcal{T}$ is the observed time period. Here $Y_{i}(\cdot)$ and $\bm{X}_{i}(\cdot)$ are considered the functional response and functional covariate for the $i$th subject, respectively, with $n$ and $p$ being the sample size and dimension of the covariates. Without loss of generality, we assume that $\mathcal{T}=[0,1]$ in this article.
@@
We focus on the relation between $Y_{i}(t)$ and $\bm{X}_{i}(t)$ for all $t\in[0,1]$, referred to as longitudinal associations in what follows. To cluster the longitudinal associations of different counties, we assume the samples $\{Y_{i}(\cdot),\bm{X}_{i}(\cdot);i=1,\ldots,n\}$ can be divided into $K$ clusters, following a finite mixture of functional linear concurrent models (FMFLCM):
$Y_{i}(t)=\sum_{j=1}^{p}X_{ij}(t)\beta_{jk}(t)+\varepsilon_{ik}(t),\quad t\in[0,1]\text{ if }Z_{i}=k,$ (1)
where $Z_{i}$ is the cluster membership for the $i$th subject, $\beta_{jk}(\cdot)$ is the functional coefficient for the $k$th group reflecting the longitudinal association, and $\{\{\varepsilon_{ik}(t);t\in[0,1]\};i=1,\ldots,n,k=1,\ldots,K\}$ are mean-zero Gaussian processes on $[0,1]$ that are independent across different $i$ and $k$. We assume that $Z_{1},\ldots,Z_{n}$ are independent and identically distributed (i.i.d.) samples from a multinomial distribution on ${1,\ldots,K}$ with $\mathbb{P}(Z_{i}=k)=\pi_{k}$ and $\operatorname{var}(\varepsilon_{ik}(t))=\sigma_{k}^{2}$, $i=1,\ldots,n$ and $t\in[0,1]$, where the error process $\{\varepsilon_{ik}(t);t\in[0,1]\}$ may possess temporal correlations among different time $t$. By taking a fixed time $t\in\mathcal{T}$ for the functional objects in (1), this model is essentially a finite mixture of linear regression models (Peel and MacLahlan (2000); Khalili and Chen (2007); Khalili and Lin (2013)), capturing heterogeneity in the associations between $\{Y_{i}(t);i=1,\ldots,n\}$ and $\{(X_{i1}(t),\ldots,X_{ip}(t))^{\top};i=1,\ldots,n\}$ for each $t$. We generalize the mixture of linear regression models to its functional version in (1), aiming to account for the heterogeneity and smoothness of $\beta_{jk}(\cdot)$ during model estimation.
@@
In the following, let $\bm{\beta}(t)=\{\beta_{jk}(t);j=1,\ldots,p;k=1,\ldots,K\}$, $\bm{\sigma}^{2}:=(\sigma_{1}^{2},\ldots,\sigma_{K}^{2})^{\top}$, and $\bm{\pi}:=(\pi_{1},\ldots,\pi_{K})^{\top}$. We use $\bm{\beta}_{\cdot k}(t)$ and $\bm{\beta}_{j\cdot}(t)$ to denote the $k$th column and the $j$th row of $\bm{\beta}(t)$, respectively. For ease of notation, we might use $\bm{\beta}$, $\bm{\beta}_{\cdot k}$, $\bm{\beta}_{j\cdot}$, and $\beta_{jk}$ to represent $\bm{\beta}(\cdot)$, $\bm{\beta}_{\cdot k}(\cdot)$, $\bm{\beta}_{j\cdot}(\cdot)$, and $\beta_{jk}(\cdot)$, respectively.
We assume the functional variables $Y_{i}(\cdot)$ and $\bm{X}_{i}(\cdot)$ in FMFLCM (1) are only observed on a finite time grid, denoted by $\{(Y_{i}(t_{is}),\bm{X}_{i}(t_{is}))$; $i=1,\ldots,n,s=1,\ldots,S_{i}\}$, where $S_{i}$ is the observed number of time points for subject $i$, and $\{t_{is};i=1,\ldots,n,s=1,\ldots,S_{i}\}$ are time points contained in $[0,1]$ that can be varying across different subjects. Based on the above notations, we estimate the parameters $\bm{\Phi}:=(\bm{\beta},\bm{\sigma}^{2},\bm{\pi})$ by maximizing the composite log-likelihood of the data (Varin et al. (2011))
$l(\bm{\Phi})$ $=\sum_{i=1}^{n}\sum_{s=1}^{S_{i}}\operatorname{log}\Biggl{\{}\sum_{k=1}^{K}\pi_{k}f\bigl{(}Y_{i}(t_{is});\bm{X}_{i}(t_{is})^{\top}\bm{\beta}_{\cdot k}(t_{is}),\sigma_{k}^{2}\bigr{)}\Biggr{\}},$ (2)
where $f(\cdot;\mu,\sigma^{2})$ is a Gaussian density function with mean $\mu$ and variance $\sigma^{2}$. In the composite likelihood, we do not consider the temporal dependencies within residual processes $\varepsilon_{ik}(\cdot)$ in order to reduce computational costs required for model estimation; similar strategies have been adopted in the estimation of functional regressions (Huang et al. (2002); Zhu et al. (2012)).
In general, optimizing $\bm{\beta}$ in (2) is an ill-posed problem due to the infinite-dimensional nature of $\bm{\beta}$. Furthermore, estimating $\bm{\beta}$ may suffer from the curse of dimensionality when $p$ is relatively large. In such cases the estimated functional coefficients can be far from their true values or may not even exist (Wang et al. (2008); Yi and Caramanis (2015)). To address these issues, we introduce a class of penalties in Section 2.3 to regularize the smoothness and sparsity of $\bm{\beta}$ simultaneously. Based on these penalties, we develop a regularized expectation maximization algorithm in Section 2.4 to simultaneously perform clustering of subjects, estimation of $\bm{\beta}$, and variable selection of the covariates.
@@ 2.3 Sparsity and Smoothness Pursued Penalties
In this subsection we propose a class of smoothly clipped absolute deviation (SCAD) type penalties (Fan and Li (2001)) to adjust the aforementioned ill-posed problems. The SCAD penalty takes the form
$\bm{P}_{\text{SCAD}}(u;\lambda)=\begin{cases}\lambda u&\text{if }0\leq u\leq\lambda,\\ -\frac{u^{2}-2\gamma\lambda u+\lambda^{2}}{2(\gamma-1)}&\text{if }\lambda<u<\gamma\lambda,\\ \frac{(\gamma+1)\lambda^{2}}{2}&\text{if }u\geq\gamma\lambda,\end{cases}$
@@
where $u$ is a scalar parameter to be penalized, $\lambda$ is a tuning parameter, and $\gamma$ is a hyperparameter chosen to be 3.7, as suggested in Fan and Li (2001). The SCAD penalty has been shown to possess desirable theoretical properties for variable selection in various regression models (Fan and Li (2001); Wang et al. (2007, 2008)). Here we extend the penalty to the functional objects $\bm{\beta}$, leading to the functional SCAD (FS) penalty
@@
$\bm{P}_{\text{FS}}(\bm{\beta};{\lambda_{1},\ldots,\lambda_{K}},r):=\sum_{j=1}^{p}\sum_{k=1}^{K}\bm{P}_{\text{SCAD}}\bigl{(}\lVert\beta_{jk}\rVert_{r};{\lambda_{k}}\bigr{)},$ (3)
@@
where $\|\beta_{jk}\|_{r}=\sqrt{\|\beta_{jk}\|^{2}+r\|\beta_{jk}^{(2)}\|^{2}}$, with $\|f\|=\sqrt{\int_{0}^{1}f^{2}(t)\,\mathrm{d}t}$ being the $L_{2}$ norm for a function $f$. The norm $\|\beta_{jk}\|_{r}$ simultaneously measures the magnitude and smoothness of $\beta_{jk}$ (Meier et al. (2009)), in which $r$ leverages the two terms in the norm. By penalizing $\bm{\beta}$ using (3), we can regularize the smoothness of each $\beta_{jk}$ to deal with its functional structure, and shrink the $\beta_{jk}$ with small norm $\|\beta_{jk}\|_{r}$ to zero (Huang et al. (2009)) for the purpose of variable selection.
@@
Recall that our SDOH functional covariates exhibit significant collinearity, as illustrated in Figure 1. This collinearity may increase variance and potentially lead to misspecification in variable selection (Zou and Hastie (2005)). To address this issue, we adopt a similar strategy as the elastic net (Zou and Zhang (2009)) to add an $L_{2}$ term to the FS penalty (3). As such, we obtain the functional SCAD-Net (FS-Net) penalty
$\bm{P}_{\text{FS-Net}}(\bm{\beta};\lambda_{1},\ldots,\lambda_{K},\rho,r):=\sum_{j=1}^{p}\sum_{k=1}^{K}\bigl{\{}\rho\bm{P}_{\text{SCAD}}\bigl{(}\lVert\beta_{jk}\rVert_{r};\lambda_{k}\bigr{)}+(1-\rho)\lambda_{k}\lVert\beta_{jk}\rVert_{r}^{2}\bigr{\}},$ (4)
@@
where $\rho\in[0,1]$ is a tuning parameter controlling the proportions between the FS penalty and the $L_{2}$ term. The above penalty not only maintains the properties of the FS penalty but also reduces variance and addresses misspecification in variable selections by penalizing the magnitude of $\sum_{j=1}^{p}\sum_{k=1}^{K}\lambda_{k}\|\beta_{jk}\|_{r}^{2}$.
@@
In general, variable selection with FS or FS-Net does not guarantee that the selected covariates from different clusters are the same. This may not be suitable for some applications where identifying important variables for all clusters is necessary. In our case, the purpose of variable selection for SDOH is to aid clustering based on the selected covariates. Therefore, we require that the selected SDOH for all clusters be the same within estimation procedures. To facilitate this kind of investigation, we propose shrinking ${\beta_{jk};k=1,\ldots,K}$ to zero functions simultaneously, referred to as cluster-invariant sparsity for variable selection. For this we modify the FS-Net penalty into the functional group SCAD-Net (FGS-Net) penalty
@@
$\bm{P}_{\text{FGS-Net}}(\bm{\beta};\lambda,\rho,r):=\sum_{j=1}^{p}\bigl{\{}\rho\bm{P}_{\text{SCAD}}\bigl{(}\lVert\bm{\beta}_{j\cdot}\rVert_{r};\lambda\bigr{)}+(1-\rho)\lambda\lVert\bm{\beta}_{j\cdot}\rVert_{r}^{2}\bigr{\}},$ (5)
@@
where $\|\bm{\beta}_{j\cdot}\|_{r}=\sqrt{\sum_{k=1}^{K}\|\beta_{jk}\|_{r}^{2}}$. By grouping $\{\beta_{kj};k=1,\ldots,K\}$ for each $j$ together in the norm, the FGS-Net not only yields cluster-invariant sparsity but also improves statistical efficiency for variable selection by borrowing strength across all clusters.
@@ 2.4 Regularized EM Algorithm
In this subsection we apply an EM algorithm to proceed with the parameter estimation of FMFLCM in (1). We first introduce a latent variable $z_{ik}:=I(Z_{i}=k)$ to represent the cluster membership of the subject $i$, where $I(\cdot)$ is an indicator function. Denote $\bm{Z}$ as $\{z_{ik};i=1,\ldots,n,k=1,\ldots,K\}$. The composite log-likelihood of FMFLCM containing $\bm{Z}$ is then given by
$l_{C}(\bm{\Phi},\bm{Z})$ $=\sum_{i=1}^{n}\sum_{k=1}^{K}z_{ik}\sum_{s=1}^{S_{i}}\bigl{[}\operatorname{log}\pi_{k}+\operatorname{log}\bigl{\{}f\bigl{(}Y_{i}(t_{is});\bigl{\{}\bm{X}_{i}(t_{is})\bigr{\}}^{\top}\bm{\beta}_{\cdot k}(t_{is}),\sigma_{k}^{2}\bigr{)}\bigr{\}}\bigr{]}.$ (6)
To address the ill-posed issue for estimating functional coefficients $\bm{\beta}$, we develop a Regularized EM (REM) algorithm by incorporating penalty functions proposed in Section 2.3. In the following we demonstrate the E-step and M-step of the REM algorithm equipped with the FGS-Net penalty (5), while those equipped with FS penalty (3) or FS-Net penalty (4) can be obtained similarly.
2.4.1 Procedures of REM algorithm
Let $\bm{\Phi}^{(m-1)}:=(\bm{\beta}^{(m-1)},(\bm{\sigma}^{2})^{(m-1)},\bm{\pi}^{(m-1)})$ be the parameters updated at the $(m-1)$th iteration. The REM algorithm iterates between the following E-step and M-step:
E-step: With the parameter $\bm{\Phi}^{(m-1)}$, we first compute the conditional expectation of $z_{ik}$, given the samples $\bm{Y}=\{Y_{i}(t_{is});i=1,\ldots,n,s=1,\ldots.S_{i}\}$ and $\bm{X}=\{\bm{X}_{i}(t_{is});i=1,\ldots,n,s=1,\ldots,S_{i}\}$,
$\begin{split}\omega_{ik}^{(m)}&:=\mathbb{E}_{\bm{\Phi}^{(m-1)}}(z_{ik}|\bm{Y},\bm{X})\\ &=\frac{{\pi_{k}^{(m-1)}}\sum_{s=1}^{S_{i}}f(Y_{i}(t_{is}),\bm{X}_{i}(t_{is})^{\top}\bm{\beta}_{\cdot k}^{(m-1)}(t_{is}),(\sigma_{k}^{2})^{(m-1)})}{\sum_{k=1}^{K}{\pi_{k}^{(m-1)}}\sum_{s=1}^{S_{i}}f(Y_{i}(t_{is}),\bm{X}_{i}(t_{is})^{\top}\bm{\beta}_{\cdot k}^{(m-1)}(t_{is}),(\sigma_{k}^{2})^{(m-1)})}.\end{split}$ (7)
After that, we calculate the expectation of the log-likelihood in (6) conditioning on $\bm{Y}$ and $\bm{X}$, given the parameter $\bm{\Phi}^{(m-1)}$. This leads to the Q function
$Q\bigl{(}\bm{\Phi}|\bm{\Phi}^{(m-1)}\bigr{)}:=\sum_{i=1}^{n}\sum_{k=1}^{K}\omega_{ik}^{(m)}\sum_{s=1}^{S_{i}}\bigl{[}\operatorname{log}\pi_{k}+\operatorname{log}\bigl{\{}f\bigl{(}Y_{i}(t_{is}),\bigl{\{}\bm{X}_{i}(t_{is})\bigr{\}}^{\top}\bm{\beta}_{\cdot k}(t_{is}),\sigma_{k}^{2}\bigr{)}\bigr{\}}\bigr{]}.$ (8)
M-step: To facilitate the update of $\bm{\beta}$, we incorporate the FGS-Net penalty (5) into the Q function (8),
$Q^{\text{pen}}\bigl{(}\bm{\Phi}|\bm{\Phi}^{(m-1)};\lambda,\rho,r\bigr{)}:=Q\bigl{(}\bm{\Phi}|\bm{\Phi}^{(m-1)}\bigr{)}-\bm{P}_{\text{FGS-Net}}(\bm{\beta};\lambda,\rho,r).$ (9)